CONTRIBUTED TALKS

Ahn, Eunkyung: Thursday, June 19, 17:45, Room 4

An extended Lie-Trotter formula and its applications (with Sejong Kim and Yongdo Lim)

Al Zhour, Zeyad: Thursday, June 19, 11:25, Room 5

Matrix Results on Weighted Drazin Inverse and Some Applications (with Adem Kilicman)

Andjelic, Milica: Friday, June 20, 10:35, Room 3

One upper bound for the largest eigenvalue of the signless Laplacian (with Slobodan Simic)

Arav, Marina: Tuesday, June 17, 10:35, Room 3

Sign Patterns That Require Almost Unique Rank (with Frank Hall, Zhongshan Li, Assefa Merid, Yubin Gao)

Barría, José: Friday, June 20, 16:20, Room 4 The strong closure of the similarity orbit for a class of pairs of finite rank operators

Bengochea, Gabriel: Friday, June 20, 11:00, Room 4 Duality in the Hopf Algebra of multivariate polynomials

Bourgeois, Gerald: Monday, June 16, 17:20, Room 4 About the logarithm function over the matrices

Bru, Rafael: Monday, June 16, 18:10, Room 3

On some classes of H-matrices (with Ljiljana Cvetković, Vladimir Kostić and Francisco Pedroche)

Bueno Cachadina, María Isabel: Tuesday, June 17, 17:20, Room 4 Algorithms for computing the Geronimus Transformation

Carriegos, Miguel: Tuesday, June 17, 11:50, Room 4 Reachability of regular switched linear systems

Castro-González, Nieves: Friday, June 20, 11:25, Room 4

Representations for the generalized Drazin inverse of additive perturbations (with M.F. Martínez-Serrano)

Cheng, Wei: Friday, June 20, 17:10, Room 4

One Type of Inverse Eigenvalue Problems in Quaternionic Quantum Mechanics (with Liang-gui Feng)

Corral, Cristina: Monday, June 16, 11:10, Room 3 On Schur complements of H-matrices (with R. Bru, I. Giménez, J. Mas)

Cortés, Vanesa: Friday, June 20, 15:30, Room 3 Some properties of the class sign regular matrices and its subclasses (with J.M. Peña)

Costa, Liliana: Monday, June 16, 10:45, Room 4 Acyclic Birkhoff Polytope (with C.M. da Fonseca and Enide Andrade Martins)

Cravo, Glória: Tuesday, June 17, 10:35, Room 4 Controllability of Matrices with Prescribed Blocks

da Cruz, Henrique F.: Thursday, June 19, 18:10, Room 4

On the matrices that preserve the value of the immanant of the upper triangular matrices (with Rosário Fernandes)

Day, Jane: Thursday, June 19, 16:55, Room 3 Graph Energy Change Due to Edge Deletion

Sánchez Perales, Salvador: Thursday, June 19, 11:00, Room 3 Manifold of proper elements (with S. Djordevic)

Dodig, Marija: Monday, June 16, 10:45, Room 3 Singular systems, state feedback problem

Dolinar, Gregor: Thursday, June 19, 17:20, Room 4 General preservers of quasi-commutativity (with Bojan Kuzma)

Domínguez, María Elena: Monday, June 16, 17:20, Room 3 General solution of certain matrix equations arising in filter design applications

Dopazo, Esther: Friday, June 20, 10:35, Room 4

Further results on the representation of the Drazin inverse of a 2 × 2 block matrix (with M.F. Martínez-Serrano and N. Castro-González)

Esen, Özlem: Friday, June 20, 17:10, Room 3 On The Root Clustering of Matrices

Fernandes, Rosário: Monday, June 16, 11:10, Room 4 Rank partitions and covering numbers under small perturbations of an element

Ferrer, Josep: Tuesday, June 17, 12:15, Room 4

Geometric structure of the equivalence classes of a controllable pair (with A. Compta, and M. Peña)

Fonseca, Carlos: Thursday, June 19, 12:15, Room 3 An inequality for the multiplicity of an eigenvalue

Fosner, Ajda: Monday, June 16, 11:35, Room 3 Commutativity preserving maps on real matrices

Frank, Martin: Tuesday, June 17, 17:45, Room 4 An iterative method for transport equations in radiotherapy (with Bruno Dubroca)

Furtado, Susana: Thursday, June 19, 11:25, Room 3 Order Invariant Spectral Properties for Several Matrices (with Charles Johnson)

Furuichi, Shigeru: Monday, June 16, 16:55, Room 4 On trace inequalities for products of matrices Gassó, Maria T.: Tuesday, June 17, 12:15, Room 3

The class of Inverse-Positive matrices with checkerboard pattern (with Manuel F. Abad, and Juan R. Torregrosa)

Glebsky, Lev: Tuesday, June 17, 11:50, Room 3 On low rank perturbations of matrices (with Luis Manuel Rivera)

Goldberger, Assaf: Friday, June 20, 16:20, Room 3

An upper bound on the characteristic polynomial of a nonnegative matrix leading to a proof of the Boyle–Handleman conjecture (with Michael Neumann)

Gouveia, María: Monday, June 16, 16:55, Room 3 On a singular Toeplitz pencil

Grudsky, Sergey: Monday, June 16, 18:35, Room 3 Uniform boundedness of Toeplitz Matrices with variable coefficients

Guzmán, José Ramón: Friday, June 20, 15:30, Room 4

Reduction of an Ito’s diffusion input output model for the determination of square mean stability

Hanaish, Ibrahim: Thursday, June 19, 17:45, Room 5

Shrinkage Estimators for Estimation the Multivariate Normal Mean Vector under Degrees of Distrust (with Abdunnabi M. Ali Elbouzedi)

Horn, Roger: Monday, June 16, 12:00, Room 3

A Canonical Form for Quasi-Real Normal Matrices Under Real Orthogonal Similarity (with Geoffrey R. Goodson)

Im, Bokhee: Thursday, June 19, 11:00, Room 5

Representations of trilinear products in Comtrans algebras (with Jonathan D. H. Smith)

Kaibah, Hussein: Thursday, June 19, 18:10, Room 5

Asymptotic Behavior of Solutions of Stochastic Equations and Applications in Statistical Parameter Estimation

Kopparty, Bhaskara Rao: Monday, June 16, 18:10, Room 2 Generalized inverses of infinite matrices

Kortesi, Peter: Thursday, June 19, 10:35, Room 5 Using Linear Algebra in Teaching Hamilton Quaternions and Graphs

Lee, Gue Myung: Thursday, June 19, 18:35, Room 4

Complexity Analysis of the Primal-Dual Interior Point Method for Second-order Cone Optimization Problem (with Bo Kyung Choi)

Lee, Hosoo: Monday, June 16, 18:35, Room 4

Contractions and nonlinear matrix equations on positive definite cones (with Yongdo Lim)

Machado, Silvia: Monday, June 16, 16:55, Room 2

GPEA’s researches about the meta resources in teaching and learning the notion of basis of a vector space (with Bianchini, B.L. and Maranhão, M.C.S.A.)

Marovt, Janko: Thursday, June 19, 11:00, Room 4

Homomorphisms of matrix semigroups over division rings from dimension two to three (with Gregor Dolinar)

Marques, Maria-da-Graça: Thursday, June 19, 16:55, Room 4 Some conditions for the commutativity of matrix patterns (with C.R. Johnson)

Martínez, José-Javier: Tuesday, June 17, 18:35, Room 4 Polynomial regression in the Bernstein basis (with Ana Marco)

Martins, Enide: Friday, June 20, 11:00, Room 3

On the spectra of some graphs like weighted rooted trees (with Rosário Fernandes and Helena Gomes)

Mikkelson, Rana: Thursday, June 19, 17:20, Room 3 Minimum Rank of Graphs with Loops

Milligan, Thomas: Friday, June 20, 15:55, Room 4 On Euclidean Squared Distance Matrices (with Chi-Kwong Li and Michael Trossett)

Mitchell, Lon: Thursday, June 19, 18:10, Room 3

Orthogonal Removal of Vertices and Minimum Semidefinite Rank (with Sivaram Narayan)

Narayan, Sivaram: Friday, June 20, 12:15, Room 3 Linearly Independent Vertices and Minimum Semidefinite Rank

Olshevsky, Vadim: Thursday, June 19, 11:25, Room 4

Can One Invert a Matrix via Graph Manipulations? (with Tom Bella and Pavel Zhlobich)

Palma, Alejandro: Thursday, June 19, 11:50, Room 4

Solution of the linear time-dependent potential by using a solvable Lie algebra^{∗} (with M. Villa, and L. Sandoval)

Patricio, Pedro: Friday, June 20, 11:50, Room 4 Some additive results on Drazin Inverses (with R.E. Hartwig)

Peña, Marta: Tuesday, June 17, 11:00, Room 4

Perturbations preserving conditioned invariant subspaces (w ith A. Compta and J. Ferrer)

Perdigão, Cecília: Thursday, June 19, 17:45, Room 3 On the equivalence class graph (with Rosário Fernandes)

Poole, George: Monday, June 16, 17:20, Room 2 Whatever Happened to Rook’s Pivoting?

Prokip, Volodymyr: Thursday, June 19, 16:55, Room 5 On the problem of diagonalizability of matrices over a principal ideal domain

Protasov, Vladimir: Wednesday, June 18, 11:50, Room 4 p-radii of linear operators and equations of self-similarity

Pruneda, Rosa E.: Tuesday, June 17, 11:00, Room 3

Complete Orthogonal Decomposition Compared with Direct Projection Methods (with Beatriz Lacruz)

Rakotondrajao, Fanja: Thursday, June 19, 17:20, Room 5 Euler’s difference table and maximum permanents of (0, 1)-matrices

Roca, Alicia: Tuesday, June 17, 11:25, Room 4 Pencils with Prescribed Constant Subpencils (with F.C. Silva)

Rodríguez, Juan Alberto: Friday, June 20, 11:50, Room 3 The Laplacian Spectrum of Hypergraphs (with Aida Kamisalic)

Rump, Siegfried M.: Monday, June 16, 17:45, Room 3

The ratio between the Toeplitz and the unstructured condition number (with H. Sekigawa)

Russo, Maria Rosaria: Tuesday, June 17, 16:55, Room 4

On some general determinantal identities of Sylvester type (with Michela Redivo-Zaglia)

Salam, Ahmed: Tuesday, June 17, 18:10, Room 4 A structure-preserving Arnoldi-like method for a class of structured matrices

Satô, Kenzi: Thursday, June 19, 12:15, Room 4 The algebraic relations of curvatures of PL manifolds

Schaeffer, Elisa: Friday, June 20, 11:25, Room 3

Locally computable approximations of absorption times for graph clustering (with Pekka Orponen and Vanesa Avalos)

Schaffrin, Burkhard: Wednesday, June 18, 11:00, Room 4

Total least-squares regularization of Tykhonov type and an ancient racetrack in Corinth (with Kyle Snow)

Sebeldin, Anatoly: Thursday, June 19, 11:50, Room 5

Algorithm resolving problem of determination of finite cyclicgroup by its automorphism group (with V.K. Vildanov and A.L. Sylla)

Seddighin, Morteza: Wednesday, June 18, 12:15, Room 4 Matrix Optimization in Statistics Semrl, Peter: Tuesday, June 17, 11:25, Room 3

Locally linearly dependent operators

Sendov, Hristo: Thursday, June 19, 11:50, Room 3 Spectral Manifolds (with A. Daniilidis, J. Malick, and A. Lewis)

Shahryari, Mohammad: Thursday, June 19, 12:15, Room 5 Z_{2}-graded symmetry classes of tensors

Sinkovic, John: Thursday, June 19, 18:35, Room 3

An upper bound for the maximum nullity of a symmetric matrix whose graph is outerplanar

Sivic, Klemen: Thursday, June 19, 10:35, Room 4 On varieties of commuting triples

Soares, Graça: Monday, June 16, 17:45, Room 4 Inequalities on an indefinite inner product space (with N. Bebiano et al.)

Stosic, Marko: Monday, June 16, 12:25, Room 3 On Generalized Procrustes Problem (with João Xavier)

Stuart, Jeffrey: Monday, June 16, 11:35, Room 4 Spectrally Arbitrary Ray Patterns

Teixeira Matos, Isabel: Friday, June 20, 15:55, Room 3 A Completion Problem over the Field of Real Numbers (with Fernando C. Silva)

Uchiyama, Mitsuru: Monday, June 16, 18:10, Room 4 A New Majorization between functions

Uhlig, Frank: Wednesday, June 18, 10:35, Room 4 Convex and Non-convex Optimization Problems for the Field of Values of a Matrix

Vallejo, Ernesto: Monday, June 16, 12:25, Room 4 Additivity obstructions for integral matrices and pyramids (with Miguel Santoyo)

Vargas, Xaab Nop: Monday, June 16, 17:45, Room 2 Students difficulties with concept of vector space from point of view of APOS Theory

Verde-Star, Luis: Friday, June 20, 12:15, Room 4 Linear algebraic approach to rational functions

Vieira, Luis: Monday, June 16, 18:35, Room 2

Euclidean Jordan algebras and inequalities on the parameters and on the spectra of a strongly regular graph

Weaver, James: Monday, June 16, 12:00, Room 4 Nonsingularity of Divisor Tournaments (with Rohan Hemasinha and Jeffrey L. Stuart)

Wojciechowski, Piotr: Friday, June 20, 16:45, Room 3 Orderings of matrix algebras and their applications Wrobel, Iwona: Thursday, June 19, 10:35, Room 3

The Gauss-Lucas theorem and the numerical range

Wu, Pei Yuan: Wednesday, June 18, 11:25, Room 4 Numerical ranges of nilpotent operators (with Hwa-Long Gau)